OpenStudy (anonymous):

How many yards of material from a 24-yard length of cloth remain after 3 pieces, each 3 yards long, and 5 pieces, each 2 yards long, are removed?

OpenStudy (anonymous):

5 yards remain. You start out with 24 yards, then you take away 3 pieces each 3 yard long, or 3*3=9. Then, you take away 5 pieces, each 2 yards long, or 5*2=10. 24-9-10=5